# Investigating the Representation of Backchannels and Fillers in Fine-tuned Language Models

### v1

@@ Abstract
Backchannels and fillers are important linguistic expressions in dialogue, but are under-represented in modern transformer-based language models (LMs). Our work studies the representation of them in language models using three fine-tuning strategies. The models are trained on three dialogue corpora in English and Japanese, where backchannels and fillers are preserved and annotated, to investigate how fine-tuning can help LMs learn their representations. We first apply clustering analysis to the learnt representation of backchannels and fillers, and have found increased silhouette scores in representations from fine-tuned models, which suggests that fine-tuning enables LMs to distinguish the nuanced semantic variation in different backchannel and filler use. We also use natural language generation (NLG) metrics to confirm that the utterances generated by fine-tuned language models resemble human-produced utterances more closely. Our findings suggest the potentials of transforming general LMs into conversational LMs that are more capable of producing human-like languages adequately.
@@ 1 Introduction
In our daily life, it is quite ubiquitous to encounter conversation such as below:
Alice: Did you check your inbox?
Bob: uh-huh.
Alice: Good.
Bob: uh, but why?
@@
The bold words in this dialogue are examples for backchannels (‘uh-huh’) and fillers (‘uh’), which play an important role in managing the flow of conversation. The use of backchannels/fillers is seen as an important way for interlocutors to negotiate the common ground (Clark, 1996). Since backchannels/fillers usually do not convey information directly and have only pragmatic functions Jucker and Smith (1998) (e.g., expressing affirmation or disagreement to the previous utterance), they are semantically ‘bleached’ elements (Fuller, 2003) and deemed optional as they lack concrete meanings (Schourup, 1999; Fuller, 2003). As a result, the NLP community often treats them similarly to ‘stop words’, excluding them during pre-processing as a method to clean the data and improve accuracy (e.g., see Sarica and Luo, 2021). For instance, in dependency parsing for spoken dialogue data, studies such as Charniak and Johnson (2001), Jørgensen (2007) and Dobrovoljc and Martinc (2018) report that excluding backchannels/fillers such as ‘uh’ and ‘er’ in the Switchboard dataset (Holliman et al., 1992) can significantly improve parsing accuracy.
@@
Although this has been a common practice, previous studies have highlighted that backchannels and fillers can contain rich contextual meaning in dialogue: backchannels such as ‘yeah’, ‘okay’, for example, can serve as feedback towards previous speech (Clark, 1996, pp. 32). Fillers, such as ‘uh’, ‘um’, are used as signals of disfluency (Rose, 2015) and reflect the speaker’s cognitive processing when searching for the next word (Clark and Fox Tree, 2002), thus indicating the speaker’s cognitive load (e.g., see Table 1 of Berthold and Jameson, 1999 and Rose, 2015). Thus, fillers can be regarded as an important signal to the listener that the speaker needs some time to complete the utterance (and wants to hold the turn; Ball, 1975). Further studies using qualitative approaches show that backchannels/fillers, as a source of feedback, play an important role in the incremental updating of dialogue from a semantic perspective, controlling the flow of information during conversation (Bergey and DeDeo, 2024) and improving the mutual understanding between the interlocutors in order to reach a joint goal (Hough et al., 2015; Howes and Eshghi, 2021).
@@
In order to figure out the potential of transforming language models into conversational language models, which can utilise backchannels/fillers, it is essential to establish how effectively they can be learnt and represented in language models (LMs). The obstacle for answering the question is that, as emphasized in Liesenfeld and Dingemanse (2022) and Dingemanse and Liesenfeld (2022), most language data in NLP is text-based. Well-formatted and well-annotated interaction data such as conversation data, is often insufficient in size and quality. Consequently, there is limited contextual information on backchannels/fillers for LMs to learn from, and thus LM-based automatic speech recognition (ASR) systems will perform poorly in recognise turn taking/holding, which is largely moderated by backchannels and fillers. Moreover, due to the absence of backchannels and fillers during the pre-training phase, most pre-trained LMs have limited knowledge of these elements. Consequently, the content generated by these LMs tends to be text-like and distant from real dialogue in its form [^1] In our opinion, this case is also applicable to LLM such as LLaMA-3-8B, as evidenced in our observation shown in the examples in Figure 6(a) to 6(d) in the Appendix, during the NLG task, the pre-trained LLM barely generates any backchannels and fillers. , which, as result, fails to ‘[put] natural in natural language processing’ (Chrupała, 2023) and lack the ability to act as competent dialogue agents that can, for example, produce (i) appropriate backchannels as feedback to user utterances, and are therefore not considered attentive enough to meet human user expectations (Buschmeier and Kopp, 2018), and (ii) natural fillers in their utterances, which are vital for their role in organising speech and their communicative functions in spoken language understanding (Dinkar et al., 2022). As a result, an LM which cannot use backchannels/fillers properly cannot be used for building a competent conversational agent.
@@
Our work presented here therefore addresses backchannels/fillers as important linguistic phenomena that are not well reflected in language models and specifically investigates the issue of learning the representations of backchannels/fillers in language models. A solution that proved effective in tackling dialogue phenomena issue, is fine-tuning (see, e.g., Noble and Maraev, 2021). Although there has been a great deal of work attempting to answer the question of how the representation of linguistic knowledge by a language model, such as BERT (Devlin et al., 2019) and GPT2 Radford et al. (2019), is altered after fine-tuning, we believe our study makes a meaningful contribution, as an initial study trying to answer how and how well language models can represent backchannels/fillers after fine-tuning. Our general research question (RQ1) and three more specific research questions are:
@@
Can modern transformer-based language models, such as BERT, GPT-2 and a larger language model such as LLaMA-3 8B, Qwen 3 8B learn representations for backchannels/fillers through fine-tuning?
@@
Do different fine-tuning strategies make a difference in learning the representation of backchannels/fillers?
2 Background
2.1 Backchannels/Fillers as Discourse Markers
Backchannels/fillers are considered to be discourse marker Jucker and Smith (1998); Jucker (1998). They are semantically ‘bleached’ elements in conversation, such as oh, yeah, uh, uhm which can be either fillers or backchannels – depending on whether they are within the dialogue as a sign of disfluency, or stand alone or at the beginning of an utterance, which are then usually taken as feedback to the previous utterance.
@@ 2.1 Backchannels/Fillers as Discourse Markers
2.2 Fine-tuning Language Models
Fine-tuning is an important step to adapt a pre-trained model to novel downstream tasks and learn representations that are important for downstream tasks. It has been consistently reported that fine-tuning can improve LMs’ representation capabilities at different levels of linguistic representation. Mosbach et al. (2020), for example, use three sentence level classification tasks selected from the GLUE benchmark (Wang et al., 2018) as the fine-tuning tasks on BERT Devlin et al. (2019), and find that fine-tuning can indeed affect the representation of linguistic knowledge in language models, especially the last hidden layers. It has been approved in many previous studies that fine-tuning is indispensable for an LM to perform well in different downstream tasks (e.g., Noble and Maraev, 2021; Merchant et al., 2020). A classical way to evaluate a fine-tuned LM is to use probing techniques to investigate the meaning representation of the hidden layer weights in the fine-tuned model, which can either be a supervised method (building a classifier to report accuracy), or an unsupervised method (using clustering to report clustering quality before and after fine-tuning) (see, e.g., Zhou and Srikumar, 2021; Mosbach et al., 2020). However, what is unknown is how language inputs that are less represented in the original language model would be represented in the fine-tuned language model [^2] We observed that backchannels/fillers have high token IDs, which indicates that they are not included in the original vocabulary of the tokeniser, or the tokenizer encountered the backchannels/fillers only in a late phase of pretraining. . We will try to answer this in our case study of backchannels/fillers, mainly through the use of clustering after the fine-tuning process.
@@ 3.1 Data Selection
In order to find datasets suitable for this study, we focussed our search specifically on datasets of transcribed spoken dialogue where fillers and backchannels are properly annotated. In order to lower the complexity of the task, the number of interlocutors in the dialogues were be limited to two. We identified the following three datasets which meet our needs: Switchboard (Holliman et al., 1992) and MapTask (Anderson et al., 1991), both English; and the BTSJ 1000 Person Japanese Natural Conversation Corpus (Usami, 2023), which is in Japanese. The two English datasets together have the same size as the Japanese dataset (about 150 000150\,000 utterances).
@@ 3.2 Task definition
Figure 1: A two-person dialogue example which shows turn-taking and contains various backchannels/fillers. The uh from the first speaker is a filler while the oh, uh hmm, and tja from the second speaker are backchannels. We merge the utterances from both speaker (bottom) specifically for the NTP and Masking fine-tuning tasks: We take the utterances from the two speaker as a whole entity and combined them into one bigger sequence which reflects our consideration that the utterances from one speaker is dependent on the utterance from the other speaker while the speaker information is retained. We add the speaker IDs (e.g., <s1>) as well to let the LM know that the utterances are from two different sources. For the TTP fine-tuning task, merging is not required. We use the TurnGPT model, which will take the utterances from both speakers in a linear-time order as input and predict which words have higher turn-taking probability.
In order to learn the representation of backchannels/fillers in language models, we need to fine-tune existing language models that do not have (or have limited) knowledge of backchannels/fillers using downstream tasks where the models can learn the contextual information of these linguistic items. We use and compare three different fine-tuning tasks: masking (MASK), next token prediction (NTP) and turn taking prediction (TTP).
Figure 2: We select the backchannel ‘uh huh’ as an example to show how the embedding is obtained from language models. The pipeline consists of the following three steps: (1) Corpus of utterances of varied length, which contain the backchannel ‘uh-huh’ in, e.g., $T_{i},T_{j}$ position, in total $|X|$ samples; (2) Text encoding through fine-tuning to get the contextual vector representation of the backchannel ‘uh-huh’, $Q$ is the length of the backchannel tokens, $K$ is then the dimension of the selected hidden state; (3) Normalise the dimension of the contextual vector and build clusters for the contextual vector of the backchannel ‘uh-huh’.
Masking
Masked language modelling encourages models to utilize bidirectional context to build contextualized token representations. Inspired by the pre-training objective of BERT (Devlin et al., 2019) [^3] Although BERT is not an autoregressive model like GPT-2 or LLaMA, it is still considered as language model. , we adapt this method to the conversational setting by selectively masking backchannels/fillers.
Let $\mathbf{X}=\left(x_{1},x_{2},\ldots,x_{T}\right)$ be a tokenized sequence of utterances drawn from a dialogue, where $x_{t}$ is the $t^{\mathrm{th}}$ token and $T$ denotes the sequence length. We first identify all matching spans $\mathbf{S}=\{s_{1},s_{2},\ldots,s_{N}\}$, where each $s_{n}$ is a continuous token subsequence of backchannels or fillers, and $N$ is the total number of such spans. For each span $s_{n}\in\mathbf{S}$, define $L=\text{len}(s_{n})$, we randomly apply a masking operation $\mathbf{M}$ as follows:
@@
where $P$ denotes the probability of each operation. [^4] For the parameter value $P$, we use the default value shown in (Devlin et al., 2019) The resulting corrupted sequence $X^{\prime}$ is then encoded into contextualized representations $H=(h_{1},h_{2},\ldots,h_{T})$. For each masked span $s_{n}\in{S}$, the model predicts the original tokens through ($W$ being the classification layer weights):
@@
By learning to predict masked backchannels or fillers from contextual discourse, the model can build a better representations. As shown in Figure 1, in order to reflect the notion that backchannels/fillers are no less different from words with substantial meaning that are conditioned by their previous contexts, in practice we merge the utterances from the two sources of speakers. One of the drawbacks of doing so is that we will miss the information of the speaker of an utterance. As a solution for the fine-tuning task input, we add two special tokens to indicate to the LM the source speakers (<s1> and <s2>, as illustrated in Figure  1). This setting also applied to the following two fine-tuning strategies.
Next-token Prediction
We consider the general task as a language modelling task, i.e., estimating the probability of the next token given the previous input.
Let $\mathbf{X}=\left(x_{1},x_{2},\ldots,x_{T}\right)$ represent a sequence of tokens, where $x_{t}$ is the token at time step $t$ (either a filler/backchannel or a regular word with substantial meaning) and $T$ is the length of the sequence. A pre-trained language model $f(\cdot)$ is fine-tuned to predict the probability of the next token $x_{t+1}$ given all previous tokens $x_{1},x_{2},\dots,x_{t}$. Thus the probability of the next token $x_{t+1}$ given the previous tokens $x_{1},x_{2},\dots,x_{t}$ is expressed as:
@@
Figure 3: The t-SNE plots of English and Japanese backchannel/filler embeddings from LLaMA-3 model (one context setting, sub-figure a, c). Sub-figures (b,d) show the effect of fine-tuning (NTP) in backchannel/filler embeddings.
For this fine-tuning task, we use the Japanese Zhao and Sawada ; Sawada et al. (2024) and English Radford et al. (2019) GPT-2 models, as well as the multilingual LLaMA-3 8B AI@Meta (2024) and Qwen-3 8B Yang et al. (2025).
Turn-taking Prediction
Our third fine-tuning task is turn-taking prediction. The task is chosen due to the close correlation between the use of backchannels/fillers and turn-taking/turn-holding. Specifically, we use the framework TurnGPT (Ekstedt and Skantze, 2020), a language model based on GPT-2 and designed for the prediction of turn-shifts in spoken dialogue. A formal definition of the turn-taking prediction task is as follow:
Let $\mathbf{X}=\left(x_{1},x_{2},\dots,x_{T}\right)$ denote a sequence of tokens consisting of the linear ordering of utterances from both interlocutors (e.g., most tokens are from interlocutor A while interlocutor B produces backchannels/fillers such as ‘uh’ or ‘um’), where $x_{t}$ is the $t^{\text{th}}$ token, and $T$ is the sequence length. The fine-tuning task is then to estimate the probability distribution $P\left(y^{*}\mid X\right)$, where $y\in\{1,2,\dots,T\}$ indicates the likelihood of a turn-taking event occurring after token $x_{y}$. The final predicted turn-taking location is then based on:
@@
Similar to the NTP task, training TurnGPT also requires speaker information for each utterance. The details on our data pre-processing are summarised in Appendix B.
@@
Figure 4: Average silhouette (SIL) scores of backchannels/fillers before and after masking, and two other fine-tuning methods. NTP (red line) is applied to both GPT-2 based models, Qwen-3 8B and LLaMA-3 8B while TTP (blue line) is only used for GPT-2 based models. For BERT, fine-tuning is performed using the MASK (green line) approach. The full context setting is not available for the GPT-2 and BERT based models given the model’s input size limitations. Dashed lines are results without fine-tuning, solid lines are those obtained after fine-tuning.
3.3 Experiment Set-Up

### v2

@@ Abstract
Backchannels and fillers are important linguistic expressions in dialogue, but often treated as ‘noise’ to be bypassed in modern transformer-based language models (LMs). Here, we study how they are represented in LMs using three fine-tuning strategies on three dialogue corpora in English and Japanese, in which backchannels and fillers are both preserved and annotated. This allows us to investigate how fine-tuning can help LMs learn these representations. We first apply clustering analysis to the learnt representation of backchannels and fillers, and find increased silhouette scores in representations from fine-tuned models, which suggests that fine-tuning enables LMs to distinguish the nuanced semantic variation in different backchannel and filler use. We also employ natural language generation metrics and qualitative analyses to verify that utterances produced by fine-tuned LMs resemble those produced by humans more closely. Our findings suggest the potential for transforming general LMs into conversational LMs that can produce human-like language more adequately.
Investigating the Representation of Backchannels and Fillers
in Fine-tuned Language Models
[…]
@@ 1 Introduction
$\lozenge/\blacklozenge$ Equal contribution.
In everyday conversations, such as the one below from the Switchboard corpus (Godfrey and Holliman, 1993, 3508:7–12), backchannels (e.g., uh-huh) and fillers (e.g., uh) are ubiquitous:
A:
Well, we have saved our newspapers for years and years –
B:
Uh-huh
A:
– because the, uh, Boy Scouts –
B:
Uh-huh
A:
– our boys have been involved in have, uh, had a huge recycling bin …
@@
They play an important role in managing the flow of conversation and are an important way for interlocutors to negotiate the common ground (Clark, 1996). Since backchannels/fillers usually do not convey information directly and have only pragmatic functions (e.g., expressing affirmation or disagreement to the previous utterance; Jucker and Smith, 1998), they are considered semantically ‘bleached’ elements (Fuller, 2003) and deemed optional due to their lack concrete meanings (Schourup, 1999). As a result, the NLP community often treats them similarly to stop words, excluding them during pre-processing as a method to “clean” the data and improve accuracy (e.g., see Sarica and Luo, 2021). In dependency parsing for spoken dialogue data, for example, studies report that excluding backchannels/fillers in the Switchboard corpus (Godfrey and Holliman, 1993) can significantly improve parsing accuracy such as (Charniak and Johnson, 2001; Jørgensen, 2007; Dobrovoljc and Martinc, 2018).
@@
However, previous studies have highlighted that backchannels and fillers can contain rich contextual meaning in dialogue: backchannels such as yeah or okay, for example, serve as feedback in response to preceding speech (Clark, 1996, pp. 32). Fillers, such as uh and um, are used as signals of disfluency (Rose, 2015) and reflect the speaker’s cognitive processing when searching for the next word (Clark and Fox Tree, 2002), thus indicating the speaker’s cognitive load (e.g., see Table 1 of Berthold and Jameson, 1999 and Rose, 2015). Fillers can thus be regarded as an important signal to the listener that the speaker needs some time to complete the utterance (and wants to hold the turn; Ball, 1975). Further studies using qualitative approaches show that backchannels, as a source of feedback, play an important role in the incremental updating of dialogue from a semantic perspective, controlling the flow of information during conversation (Bergey and DeDeo, 2024) and improving the mutual understanding between the interlocutors in order to reach a joint goal (Howes and Eshghi, 2021).
@@
In order to figure out the potential of transforming language models (LMs) into conversational LMs, which can utilise backchannels and fillers, it is essential to establish how effectively they can be learnt and represented. An obstacle for answering this question is that language data used in NLP is often text-based Liesenfeld and Dingemanse (2022); Dingemanse and Liesenfeld (2022) and that well-annotated conversation data is often insufficient in size and quality. Consequently, there is limited contextual information on backchannels/fillers for LMs to learn from. This might be one reasons why LM-based automatic speech recognition (ASR) systems perform poorly in recognising turn taking/holding, a practice which is largely moderated by backchannels and fillers. Moreover, due to the absence of backchannels and fillers during the pre-training phase, most LMs have limited knowledge of these linguistic units. Consequently, the content generated by these LMs tends to be text-like and differs in form from natural dialogue [^1] In our opinion, this case is also applicable to LLM such as LLaMA-3-8B, as evidenced in our observation shown in the examples dialogues 1 to 4 in Figure 9 in the Appendix, during the NLG task, the pre-trained LLM barely generates any backchannels and fillers, which should be considered as an important sign that backchannels/fillers are not well represented in the pre-trained language model. , a reason, why NLP failed to “[put] natural in natural language processing” (Chrupała, 2023). Consequently LMs lack the ability to act as competent dialogue agents that can, for example, produce (i) appropriate backchannels as feedback to user utterances, and are therefore not considered attentive enough to meet human user expectations (Buschmeier and Kopp, 2018), and (ii) natural fillers in their utterances, which are vital for their role in organising speech and their communicative functions in spoken language understanding (Dinkar et al., 2022). As a result, an LM which ignores backchannels and fillers does not clean the data, it strips the model of the social cues necessary for natural and fluid interaction (Edlund et al., 2008), and consequently cannot be used for building a competent conversational agent.
@@
In this study, we therefore address backchannels/fillers as important linguistic phenomena that are not well reflected in language models and specifically investigate the issue of learning the representations of backchannels and fillers in LMs. Fine-tuning is an approach that has proven effective in tackling the issue of dialogue phenomena in LMs (see, e.g., Noble and Maraev, 2021). Although there has been a great deal of work attempting to answer the question of how the representation of linguistic knowledge in language models, such as BERT (Devlin et al., 2019) and GPT2 Radford et al. (2019), is altered after fine-tuning, we believe that work makes a meaningful contribution, as an initial study trying to answer how and how well LMs can represent backchannels and fillers after fine-tuning. Our general research question (RQ1) and three more specific research questions are:
@@
Does fine-tuning improve the representations of backchannels and fillers in modern transformer-based language models, such as BERT, GPT-2 and a larger language model such as LLaMA-3 8B, Qwen-3 8B?
@@
Do various fine-tuning strategies make a difference in learning the representation of backchannels/fillers?
2 Related Work
Backchannels/Fillers as Discourse Markers:
Backchannels/fillers are considered to be discourse marker Jucker and Smith (1998). They are semantically ‘bleached’ elements in conversation, such as oh, yeah, uh, uhm which can be either fillers or backchannels – depending on whether they are within the dialogue as a sign of disfluency, or stand alone or at the beginning of an utterance, which are then usually taken as feedback to the previous utterance.
@@ 2 Related Work
Fine-tuning and Representation:
Fine-tuning is an important step to adapt a pre-trained model to novel downstream tasks and learn representations that are important for downstream tasks. It has been consistently reported that fine-tuning can improve LMs’ representation capabilities at different levels of linguistic representation. Mosbach et al. (2020), for example, use three sentence level classification tasks selected from the GLUE benchmark (Wang et al., 2018) as the fine-tuning tasks on BERT Devlin et al. (2019), and find that fine-tuning can indeed affect the representation of linguistic knowledge in language models, especially the last hidden layers. It has been approved in many previous studies that fine-tuning is indispensable for an LM to perform well in different downstream tasks (e.g., Noble and Maraev, 2021; Merchant et al., 2020). A classical way to evaluate a fine-tuned LM is to use probing techniques to investigate the meaning representation of the hidden layer weights in the fine-tuned model, which can either be a supervised method (building a classifier to report accuracy), or an unsupervised method (using clustering to report clustering quality before and after fine-tuning) (see, e.g., Zhou and Srikumar, 2021; Mosbach et al., 2020). However, what is unknown is how language inputs that are less represented in the original language model would be represented in the fine-tuned language model [^2] We observed that backchannels and fillers have high token IDs, which indicates that they are not included in the original vocabulary of the tokenizer, or the tokenizer encountered the backchannels and fillers only in a late phase of pre-training. . We will try to answer this in our case study of backchannels/fillers, mainly through the use of clustering after the fine-tuning process.
@@ 3.1 Data Selection
In order to find datasets suitable for this study, we focussed our search specifically on datasets of transcribed spoken dialogue where fillers and backchannels are properly annotated. In order to lower the complexity of the task, the number of interlocutors per dialogue were be limited to two. We identified the following three datasets to meet our needs: Switchboard (Godfrey and Holliman, 1993) and MapTask (Anderson et al., 1991), both English; and the BTSJ 1000 Person Japanese Natural Conversation Corpus (Usami, 2023), which is in Japanese. Combined, the two English datasets are the same size as the Japanese dataset (about 150 000150\,000 utterances).
@@ 3.2 Task definition
In order to learn the representation of backchannels/fillers in language models, we chose to fine-tune existing language models that do not have (or have limited) knowledge of backchannels/fillers. For fine-tuning we use downstream tasks where the models can learn the contextual information of these linguistic items. We use and compare three different fine-tuning tasks: masking (MASK), next token prediction (NTP) and turn taking prediction (TTP). These are described in the following.
Figure 1: A dialogue example which shows turn-taking and contains various backchannels/fillers. The um from the speaker S1 is a filler while the oh, uh-huh, and yeah from the speaker S2 are backchannels. For the NTP and Masking fine-tuning tasks we merge the utterances from both speakers. For this, we take the utterances from the both and combined them into one larger sequence, considering that the utterances from one speaker is dependent on the utterances from the other speaker. We retain speaker information by adding speaker IDs (e.g., <s1>) to let the LM know that the utterances are from two different sources. Merging is not required for the TTP task. TurnGPT takes the utterances from both speakers in a linear-time order as input and predicts turn-taking probabilities.
Figure 2: We select the backchannel uh-huh as an example to show how embeddings are obtained from language models. The pipeline consists of the following three steps: (1) taking a corpus of utterances of various lengths, which contain the backchannel uh-huh in, e.g., position $T_{i},T_{j}$, (in total $|X|$ samples); (2) encoding the text through fine-tuning to get the contextual vector representation of the backchannel uh-huh ($Q$ is the length of the backchannel tokens, $K$ is the dimension of the selected hidden state); (3) reducing the dimension of the contextual vector and build clusters for the contextual vector of the backchannel uh-huh.
Masking (MASK)
Masked language modelling encourages models to utilize bidirectional context to build contextualized token representations. Inspired by the pre-training objective of BERT (Devlin et al., 2019) [^3] Although BERT is not an autoregressive model like GPT-2 or LLaMA, it is still considered a language model. , we adapt this method to the conversational setting by selectively masking backchannels/fillers.
Let $\mathbf{X}=\left(x_{1},x_{2},\ldots,x_{T}\right)$ be a tokenized sequence of utterances drawn from a dialogue, where $x_{t}$ is the $t^{\mathrm{th}}$ token and $T$ denotes the sequence length. We first identify all matching spans $\mathbf{S}=\{s_{1},s_{2},\ldots,s_{N}\}$, where each $s_{n}$ is a continuous token subsequence of backchannels or fillers, and $N$ is the total number of such spans. For each span $s_{n}\in\mathbf{S}$, we define $L=\text{len}(s_{n})$ and randomly apply a masking operation $\mathbf{M}$ as follows:
@@
where $P$ denotes the probability of each operation (we use the default values of Devlin et al. 2019). The resulting corrupted sequence $X^{\prime}$ is then encoded into contextualized representations $H=(h_{1},h_{2},\ldots,h_{T})$. For each masked span $s_{n}\in{S}$, the model predicts the original tokens through ($W$ being the classification layer weights):
@@
By learning to predict masked backchannels or fillers from contextual discourse, the model can build a better representation. As shown in Figure 1, in order to reflect the notion that backchannels and fillers are no less different from words with substantial meaning that are conditioned by their previous contexts, in our experiment practice we merge the utterances from the two sources of speakers. One of the drawbacks of doing so is that we will miss the information of the speaker of an utterance. As a solution for the fine-tuning task input, we add two special tokens to indicate to the LM the source speakers (<s1> and <s2>, as illustrated in Figure 1). This setting is also applied to the fine-tuning strategy: next token prediction.
Next-token Prediction (NTP)
We consider the general task as a language modelling task, i.e., estimating the probability of the next token (backchannel/filler or other word) given previous input.
Let $\mathbf{X}=\left(x_{1},x_{2},\ldots,x_{T}\right)$ represent a sequence of tokens, where $x_{t}$ is the token at time step $t$ (either a filler/backchannel or a regular word with substantial meaning) and $T$ is the length of the sequence. A pre-trained language model $f(\cdot)$ is fine-tuned to estimate the probability of the next token $x_{t+1}$ given all previous tokens $x_{1},x_{2},\dots,x_{t}$. Thus the probability of the next token $x_{t+1}$ given the previous tokens $x_{1},x_{2},\dots,x_{t}$ is expressed as:
@@
Figure 3: t-SNE plots of the English and Japanese backchannel/filler embeddings from the LLaMA-3 model (a, c) before fine-tuning, and (b, d) after NTP fine-tuning. Setting: one-context.
For this fine-tuning task, we use the Japanese Zhao and Sawada (2021); Sawada et al. (2024) and English Radford et al. (2019) GPT-2 models, as well as the multilingual LLaMA-3 8B AI@Meta (2024) and Qwen-3 8B Yang et al. (2025) models.
Turn-taking Prediction (TTP)
Our third fine-tuning task is turn-taking prediction. The task is chosen due to the close relation between the use of backchannels/fillers and turn-taking/turn-holding. Specifically, we use the framework TurnGPT (Ekstedt and Skantze, 2020), a language model based on GPT-2 and designed for the prediction of turn-shifts in spoken dialogue. A formal definition of the turn-taking prediction task is as follows.
Let $\mathbf{X}=\left(x_{1},x_{2},\dots,x_{T}\right)$ denote a sequence of tokens consisting of the linear ordering of utterances from both interlocutors (e.g., most tokens are from interlocutor A while interlocutor B produces backchannels/fillers, where $x_{t}$ is the $t^{\text{th}}$ token, and $T$ is the sequence length. The fine-tuning task is then to estimate the probability distribution $P\left(y^{*}\mid X\right)$, where $y\in\{1,2,\dots,T\}$ indicates the likelihood of a turn-taking event occurring after token $x_{y}$. The final predicted turn-taking location is then based on:
@@
Similar to the NTP task, training TurnGPT also requires speaker ID (e.g., <s1>) for each utterance in order to reflect turn shifts. The details on our data pre-processing are summarised in Appendix B.
@@
Figure 4: Average silhouette (SIL) scores of backchannels/fillers before and after masking, and two other fine-tuning methods. NTP (red line) is applied to GPT-2 based models, Qwen-3 8B and LLaMA-3 8B, while TTP (blue line) is only used for GPT-2 based models. For BERT, fine-tuning is performed using the MASK (green line) approach. The full-context setting is not available for the GPT-2 and BERT based models given the model’s input size limitations. Dashed lines are results without fine-tuning (“no-ft”), solid lines are those obtained after fine-tuning. The robustness test on the results are summarised in Table 6 in the Appendix.
3.3 Experimental Set-Up
Our experiment workflow is shown in Figure 2. The chosen LMs are fine-tuned in advance based on the three different fine-tuning tasks [^4] Note that for our analysis, when we extract the embeddings of backchannels/fillers for the pre-trained LMs without fine-tuning, the pipeline works in the same way.. For the NTP and MASK tasks, we selected 80% of the conversation data for fine-tuning and the remaining 20% for subsequent generative evaluation. Conversely, for the TTP task, all datasets were partitioned into training, validation, and testing sets exclusively for the fine-tuning process. For the evaluation of the representations of backchannels/fillers, we used all the conversation data.